# Shuffle Products, Degenerate Affine Hecke Algebras, and Quantum Toda Lattice

### v2

@@ 2.1. Shibukawa-Ueno $R$-operator
Let us consider $\mathrm{Sym}_{k}$ and $\mathrm{Sym}_{k}^{-}$ as elements of $\mathbf{C}[S_{k}]$ embedded into semi-direct product $\mathbf{C}[S_{k}]\ltimes\mathcal{M}_{k}$. Then Proposition˜2.3 for the quantization parameter $\kappa$ is equivalent to equality
@@
Then $\varpi_{k}(\check{R}^{\theta}_{i}(\xi|-\kappa))=\check{R}^{\theta}_{i}(\xi|\kappa)$, in particular, $\varpi_{k}(\check{H}^{\theta}_{k}(-\kappa))=\check{A}^{\theta}_{k}(\kappa)$ by definition. Combining with the equality above and using $\varpi_{k}(\mathrm{Sym}_{k})=\mathrm{Sym}_{k}^{-}$, we obtain
$\check{A}^{\theta}_{k}(\kappa)=\varpi_{k}(\check{H}^{\theta}_{k}(-\kappa))=\prod_{i<j}G^{\theta}(z_{i}-z_{j},-\kappa)\mathrm{Sym}_{k}^{-},$
@@
2.2. Degenerate affine Hecke algebras and shuffle products
In this subsection, we recall some basic facts about degenerate affine Hecke algebras and relate them to Feigin-Odesskii shuffle algebras [FO98]. Unless otherwise stated, we refer the reader to [Kir97] for all the statements and proofs (upon replacing $h$ in loc. cit. with $\kappa$).
@@ 2.2. Degenerate affine Hecke algebras and shuffle products
It contains the group algebra $\mathbf{C}[S_{k}]$ of the symmetric group and the polynomial algebra $\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]$. The latter can be endowed with a $\mathfrak{H}^{\kappa}_{k}$-module structure via induction from the trivial $\mathbf{C}[S_{k}]$-representation
(2.4) $\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]\cong\mathfrak{H}^{\kappa}_{k}\otimes_{\mathbf{C}[S_{k}]}\mathbf{C}.$
@@
(2.5) $\sigma^{\kappa}_{i}=P_{i,i+1}+\kappa\frac{1-P_{i,i+1}}{Y_{i}-Y_{i+1}},$
where $P_{i,i+1}$ is the operator (2.3). It follows that the usual symmetric polynomials $\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}$ lie in the center of $\mathfrak{H}^{\kappa}_{k}$ (and, in fact, coincide with it).
By affinization procedure [KS97, Section 8.7.1], the collection of operators
@@
$\check{R}^{\mathrm{Hecke}}_{i}(\xi_{i+1}-\xi_{i+2})\check{R}^{\mathrm{Hecke}}_{i+1}(\xi_{i}-\xi_{i+2})\check{R}^{\mathrm{Hecke}}_{i}(\xi_{i}-\xi_{i+1})=$
$=\check{R}^{\mathrm{Hecke}}_{i+1}(\xi_{i}-\xi_{i+1})\check{R}^{\mathrm{Hecke}}_{i}(\xi_{i}-\xi_{i+2})\check{R}^{\mathrm{Hecke}}_{i+1}(\xi_{i+1}-\xi_{i+2}).$
@@
(2.6) $e_{k}:=\frac{1}{k!}\sum_{\sigma\in S_{k}}\sigma,\qquad e^{-}_{k}:=\frac{1}{k!}\sum_{\sigma\in S_{k}}(-1)^{\sigma}\sigma$
in $\mathbf{C}[S_{k}]$. The following is an analog of [Juc66].
@@
$e_{k}$ $=\frac{1}{k!}(\check{R}^{\mathrm{Hecke}}_{k-1}(1)\ldots\check{R}^{\mathrm{Hecke}}_{1}(k-1))\ldots(\check{R}^{\mathrm{Hecke}}_{k-1}(1)\check{R}^{\mathrm{Hecke}}_{k-2}(2))(\check{R}^{\mathrm{Hecke}}_{k-1}(1)),$
$e_{k}^{-}$ $=\frac{(-1)^{\frac{k(k+1)}{2}}}{k!}(\check{R}^{\mathrm{Hecke}}_{k-1}(-1))(\check{R}^{\mathrm{Hecke}}_{k-2}(-2)\check{R}^{\mathrm{Hecke}}_{k-1}(-1))\ldots(\check{R}^{\mathrm{Hecke}}_{1}(-k+1)\ldots\check{R}^{\mathrm{Hecke}}_{k-1}(-1))$
@@
It is enough to check the relation in some faithful representation. As such, we can take $(\mathbf{C}^{N})^{\otimes k}$ with the usual permutation action for $N\gg 0$. Then it is the content of [Mol07, Proposition 1.6.2] and [Mol07, Proposition 1.6.3]. ∎
Consider the action of Hecke $R$-matrices in the polynomial representation (2.4). Using explicit formula (2.5) and ˜2.1, we see that it is essentially given by the rational Shibukawa-Ueno braiding element:
@@
The following proposition is a rational analog of the relation between Hall-Littlewood polynomials [Mac95, Chapter III] and affine Hecke algebras (see [Lus83]), and follows immediately from Proposition˜2.6, Proposition˜2.3, and ˜2.4.
@@
The action of (anti-)symmetrizer in the representation (2.4) is given by
$(e^{\kappa}_{k}f)(Y_{1},\ldots,Y_{k})$ $=\frac{1}{k!}\mathrm{Sym}_{k}\left(\prod_{i<j}\frac{Y_{i}-Y_{j}+\kappa}{Y_{i}-Y_{j}}f(Y_{1},\ldots,Y_{k})\right),$
$((e^{-}_{k})^{\kappa}f)(Y_{1},\ldots,Y_{k})$ $=\frac{1}{k!}\prod_{i<j}\frac{Y_{i}-Y_{j}-\kappa}{Y_{i}-Y_{j}}\mathrm{Sym}_{k}^{-}f(Y_{1},\ldots,Y_{k}).$
@@
(2.7) $\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}\xrightarrow{\sim}e^{\kappa}_{k}\mathfrak{H}^{\kappa}_{k}e^{\kappa}_{k},\qquad f(Y_{1},\ldots,Y_{k})\mapsto e^{\kappa}_{k}f(Y_{1},\ldots,Y_{k})e^{\kappa}_{k}.$
@@
$\mathcal{F}^{\kappa}=\bigoplus_{k=0}^{\infty}\mathcal{F}^{\kappa}_{k},\qquad\mathcal{F}^{\kappa}_{k}=e^{\kappa}_{k}\mathfrak{H}^{\kappa}_{k}e^{\kappa}_{k}=\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}},$
@@
$(f*^{\kappa}g)(Y_{1},\ldots,Y_{k+l})=e^{\kappa}_{k+l}f(Y_{1},\ldots,Y_{k})g(Y_{k+1},\ldots,Y_{k+l})e^{\kappa}_{k+l}.$
@@
$^{N}\mathcal{F}^{\kappa}=\bigoplus_{k=0}^{\infty}{}^{N}\mathcal{F}^{\kappa}_{k},\qquad{}^{N}\mathcal{F}^{\kappa}_{k}=\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}_{<N}$
@@
$(f*g)(Y_{1},\ldots,Y_{k+l})=\frac{1}{k!l!}\mathrm{Sym}_{k+l}(f(Y_{1},\ldots,Y_{k})g(Y_{k+1},\ldots,Y_{k+l})).$
There is another deformation of the latter which is a simplest case of the Feigin-Odesskii shuffle algebras [FO98].
@@
The rational Feigin-Odesskii shuffle algebra $\mathcal{S}^{\kappa}$ is the vector space
$\mathcal{S}^{\kappa}=\bigoplus\mathcal{S}^{\kappa}_{k},\qquad\mathcal{S}^{\kappa}_{k}=\mathbf{C}[Y_{1},\ldots,Y_{k}]^{S_{k}}$
@@
$(f*^{\kappa}g)(Y_{1},\ldots,Y_{k+l})=\frac{1}{k!l!}\mathrm{Sym}_{k}\left(\prod_{\begin{subarray}{c}i=1\ldots k\\ j=k+1\ldots k+l\end{subarray}}\frac{Y_{i}-Y_{j}+\kappa}{Y_{i}-Y_{j}}f(Y_{1},\ldots,Y_{k})g(Y_{k+1},\ldots,Y_{k+l})\right)$
The $N$-truncated rational Feigin-Odesskii shuffle algebra $^{N}\mathcal{S}^{\kappa}$ is the vector space
$^{N}\mathcal{S}^{\kappa}=\bigoplus_{k=0}^{\infty}{}^{N}\mathcal{S}^{\kappa}_{k},\qquad{}^{N}\mathcal{S}^{\kappa}_{k}=\mathbf{C}[\hbar][Y_{1},\ldots,Y_{k}]^{S_{k}}_{<N}$
@@
The original construction of elliptic algebras [OF89, OF97, FO98] defines a shuffle product with factor $\vartheta_{1}(z)$ (up to shifts) on the symmetric algebra $S^{\bullet}H^{0}(\Theta_{N})$ of global sections of a degree $N$ line bundle on an elliptic curve. By [EH04], the latter admits a degeneration such that $\vartheta_{1}(z)\mapsto z$ and the symmetric algebra becomes $^{N}\mathcal{S}^{\kappa}$.
The following theorem is a corollary of Proposition˜2.7.
@@
One can similarly relate the trigonometric Shibukawa-Ueno operator to the Laurent polynomial representation of the non-degenerate affine Hecke algebra [CP95, Section 12.3] via the affinization procedure of [KS97]. It would be interesting to relate the elliptic Shibukawa-Ueno operators to elliptic affine Hecke algebras [GKV97].
2.3. Shifted Yangian of $\mathfrak{sl}_{2}$
@@
The following definition is [FPT22, Section 2.1]; we use its equivalent form in terms of currents. In particular, we assume that the weight $(m_{1},m_{2})$ is antidominant, i.e. $m_{1}\leq m_{2}$.
@@
The shifted Yangian $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})$ is an algebra over $\mathbf{C}[\hbar]$ generated by $\{e^{(i)},f^{(i)}\}_{i\geq 1}$ and $\{d_{j}^{(i)}\}_{i>m_{j}}$ for $j=1,2$ that we combine into the series
$x_{+}(u)=\sum_{i\geq 1}x_{+}^{(i)}u^{-i},\qquad x_{-}(u)=\sum_{i\geq 1}x_{-}^{(i)}u^{-i},\qquad d_{j}(u)=u^{m_{j}}+\sum_{i>m_{j}}d_{j}^{(i)}u^{-i}$
with the values in $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})[[u^{-1}]]$ satisfying
(2.8) $[d_{1}(u),d_{2}(v)]$ $=0,$
(2.9) $(u-v)[x_{+}(u),x_{-}(v)]$ $=\hbar(d_{1}(v)^{-1}d_{2}(v)-d_{1}(u)^{-1}d_{2}(u)),$
(2.10) $(u-v)[d_{j}(u),x_{+}(v)]$ $=\hbar(\delta_{j1}-\delta_{j2})d_{j}(u)(x_{+}(u)-x_{+}(v)),$
(2.11) $(u-v)[d_{j}(u),x_{-}(v)]$ $=\hbar(\delta_{j2}-\delta_{j1})(x_{-}(u)-x_{-}(v))d_{j}(u),$
(2.12) $(u-v)[x_{\pm}(u),x_{\pm}(v)]$ $=\mp\hbar(x_{\pm}(u)-x_{\pm}(v))^{2},$
The positive (negative) subalgebra $\mathrm{Y}(\mathfrak{gl}_{2})^{\pm}$ is the subalgebra of $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})$ generated by the modes of $x_{\pm}(u)$. The non-negative subalgebra $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})^{\geq 0}$ (resp. non-positive subalgebra $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})^{\leq 0}$) is the subalgebra generated by the modes of $x_{+}(u),d_{j}(u)$ (resp. $x_{-}(u),d_{j}(u)$) for $j=1,2$.
@@
$\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})^{\geq 0}\xrightarrow{\sim}\mathrm{Y}_{0,0}(\mathfrak{gl}_{2})^{\geq 0},\qquad x_{+}(u)\mapsto x_{+}(u),\ d_{j}(u)\mapsto u^{-m_{j}}d_{j}(u),$
@@
Let $m=m_{1}-m_{2}$. The definition of the shifted Yangian $\mathrm{Y}_{m}(\mathfrak{sl}_{2})$ can be found in [BFN19, Section B.1], see also its form in terms of currents $e(u),f(u),h(u)$ in [Fin+18, Section 6.1]. Its relation with $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})$ is as follows.
Proposition 2.15.
[FPT22, Proposition 2.19] There is a $\mathbf{C}[\hbar]$-algebra embedding
@@
Let $(m_{1},m_{2})=(-N,N)$. Consider the localized algebra $\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}$ of (4.5). Denote by $^{N}\tilde{\mathcal{A}}_{\hbar}:=\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}[\mathbf{u}_{1}^{\pm},\ldots,\mathbf{u}_{N}^{\pm}]$ the algebra of difference operators such that
$\mathbf{u}_{i}^{\pm}f(w)=f(w\pm\hbar\epsilon_{i})\mathbf{u}_{i}^{\pm},\ f(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}},$
where $f(w\pm\hbar\epsilon_{i}):=f(w_{1},\ldots,w_{i}\pm\hbar,\ldots,w_{N},\hbar)$. The homomorphism of the following theorem is called the GKLO homomorphism, see [Ger+05] for the non-shifted version, [Kam+14] for dominant shifts, and [BFN19] for general definition. We use the version adapted to $\mathfrak{gl}_{2}$.
Theorem 2.16.
[FPT22, Theorem 2.35] There is a $\mathbf{C}[\hbar]$-algebra homomorphism
@@
$d_{1}(u)$ $\mapsto\prod_{i=1}^{N}(u-w_{i}),\ d_{2}(u)\mapsto\prod_{i=1}^{N}(u-w_{i}-\hbar)^{-1}.$
The following is a special case of truncated shifted Yangians, see [BFN19, Appendix B]
@@
Recall the shuffle algebra of ˜2.8. The following is [Tsy21, Theorem 6.20] combined with Theorem˜2.11, see also [Enr03].
@@
3. Harish-Chandra bimodules

### v4

@@ 3.1 Shibukawa–Ueno $\boldsymbol{R}$-operator
Let us consider $\mathrm{Sym}_{k}$ and $\mathrm{Sym}_{k}^{-}$ as elements of $\mathbf{C}[S_{k}]$ embedded into semi-direct product $\mathbf{C}[S_{k}]\ltimes\mathcal{M}_{k}$. Then Proposition 3.3 for the quantization parameter $\kappa$ is equivalent to equality
@@
Then $\varpi_{k}\bigl(\check{R}^{\theta}_{i}(\xi|-\kappa)\bigr)=\check{R}^{\theta}_{i}(\xi|\kappa)$, in particular, $\varpi_{k}\bigl(\check{H}^{\theta}_{k}(-\kappa)\bigr)=\check{A}^{\theta}_{k}(\kappa)$ by definition. Combining with the equality above and using $\varpi_{k}(\mathrm{Sym}_{k})=\mathrm{Sym}_{k}^{-}$, we obtain
$\check{A}^{\theta}_{k}(\kappa)=\varpi_{k}\bigl(\check{H}^{\theta}_{k}(-\kappa)\bigr)=\prod_{i<j}G^{\theta}(z_{i}-z_{j},-\kappa)\mathrm{Sym}_{k}^{-},$
@@
3.2 Degenerate affine Hecke algebras and shuffle products
In this subsection, we recall some basic facts about degenerate affine Hecke algebras and relate them to Feigin–Odesskii shuffle algebras [12]. Unless otherwise stated, we refer the reader to [31] for all the statements and proofs (upon replacing $h$ in loc. cit. with $\kappa$).
@@ 3.2 Degenerate affine Hecke algebras and shuffle products
It contains the group algebra $\mathbf{C}[S_{k}]$ of the symmetric group and the polynomial algebra $\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]$. The latter can be endowed with a $\mathfrak{H}^{\kappa}_{k}$-module structure via induction from the trivial $\mathbf{C}[S_{k}]$-representation
$\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]\cong\mathfrak{H}^{\kappa}_{k}\otimes_{\mathbf{C}[S_{k}]}\mathbf{C}$ (3.4)
@@
$\sigma^{\kappa}_{i}=P_{i,i+1}+\kappa\frac{1-P_{i,i+1}}{Y_{i}-Y_{i+1}},$ (3.5)
where the operator $P_{i,i+1}$ is defined by (3.3). It follows that the usual symmetric polynomials $\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]^{S_{k}}$ lie in the center of $\mathfrak{H}^{\kappa}_{k}$ (and, in fact, coincide with it).
By affinization procedure [32, Section 8.7.1], the collection of operators
@@
$\check{R}^{\mathrm{Hecke}}_{i}(\xi_{i+1}-\xi_{i+2})\check{R}^{\mathrm{Hecke}}_{i+1}(\xi_{i}-\xi_{i+2})\check{R}^{\mathrm{Hecke}}_{i}(\xi_{i}-\xi_{i+1})$
$\qquad=\check{R}^{\mathrm{Hecke}}_{i+1}(\xi_{i}-\xi_{i+1})\check{R}^{\mathrm{Hecke}}_{i}(\xi_{i}-\xi_{i+2})\check{R}^{\mathrm{Hecke}}_{i+1}(\xi_{i+1}-\xi_{i+2}).$
@@
$e_{k}:=\frac{1}{k!}\sum_{\sigma\in S_{k}}\sigma,\qquad{\rm e}^{-}_{k}:=\frac{1}{k!}\sum_{\sigma\in S_{k}}(-1)^{\sigma}\sigma$ (3.6)
in $\mathbf{C}[S_{k}]$. The following is an analog of [26].
@@
$e_{k}=\frac{1}{k!}\bigl[\check{R}^{\mathrm{Hecke}}_{k-1}(1)\cdots\check{R}^{\mathrm{Hecke}}_{1}(k-1)\bigr]\cdots\bigl[\check{R}^{\mathrm{Hecke}}_{k-1}(1)\check{R}^{\mathrm{Hecke}}_{k-2}(2)\bigr]\bigl[\check{R}^{\mathrm{Hecke}}_{k-1}(1)\bigr],$
$e_{k}^{-}=\frac{(-1)^{\frac{k(k+1)}{2}}}{k!}\bigl[\check{R}^{\mathrm{Hecke}}_{k-1}(-1)\bigr]\bigl[\check{R}^{\mathrm{Hecke}}_{k-2}(-2)\check{R}^{\mathrm{Hecke}}_{k-1}(-1)\bigr]\cdots\bigl[\check{R}^{\mathrm{Hecke}}_{1}(-k+1)\cdots\check{R}^{\mathrm{Hecke}}_{k-1}(-1)\bigr].$
@@
It is enough to check the relation in some faithful representation. As such, we can take $\bigl(\mathbf{C}^{N}\bigr)^{\otimes k}$ with the usual permutation action for $N\gg 0$. Then it is the content of [38, Proposition 1.6.2] and [38, Proposition 1.6.3]. ∎
Consider the action of Hecke $R$-matrices in the polynomial representation (3.4). Using explicit formula (3.5) and Definition 3.1, we see that it is essentially given by the rational Shibukawa–Ueno braiding element
@@
The following proposition is a rational analog of the relation between Hall–Littlewood polynomials [36, Section III] and affine Hecke algebras (see [35]), and follows immediately from Proposition 3.6 and 3.3, and Corollary 3.4.
@@
The action of ((anti-))symmetrizer in the representation (3.4) is given by
$({\rm e}^{\kappa}_{k}f)(Y_{1},\dots,Y_{k})=\frac{1}{k!}\mathrm{Sym}_{k}\left(\prod_{i<j}\frac{Y_{i}-Y_{j}+\kappa}{Y_{i}-Y_{j}}f(Y_{1},\dots,Y_{k})\right),$
$(({\rm e}^{-}_{k})^{\kappa}f)(Y_{1},\dots,Y_{k})=\frac{1}{k!}\prod_{i<j}\frac{Y_{i}-Y_{j}-\kappa}{Y_{i}-Y_{j}}\mathrm{Sym}_{k}^{-}f(Y_{1},\dots,Y_{k}).$
@@
$\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]^{S_{k}}\xrightarrow{\sim}{\rm e}^{\kappa}_{k}\mathfrak{H}^{\kappa}_{k}{\rm e}^{\kappa}_{k},\qquad f(Y_{1},\dots,Y_{k})\mapsto{\rm e}^{\kappa}_{k}f(Y_{1},\dots,Y_{k}){\rm e}^{\kappa}_{k}.$ (3.7)
@@
$\mathcal{F}^{\kappa}=\bigoplus_{k=0}^{\infty}\mathcal{F}^{\kappa}_{k},\qquad\mathcal{F}^{\kappa}_{k}={\rm e}^{\kappa}_{k}\mathfrak{H}^{\kappa}_{k}{\rm e}^{\kappa}_{k}=\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]^{S_{k}},$
@@
$(f*^{\kappa}g)(Y_{1},\dots,Y_{k+l})={\rm e}^{\kappa}_{k+l}f(Y_{1},\dots,Y_{k})g(Y_{k+1},\dots,Y_{k+l}){\rm e}^{\kappa}_{k+l}.$
@@
$^{N}\!\mathcal{F}^{\kappa}=\bigoplus_{k=0}^{\infty}{}^{N}\!\mathcal{F}^{\kappa}_{k},\qquad{}^{N}\!\mathcal{F}^{\kappa}_{k}=\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]^{S_{k}}_{<N}$
@@
$(f*g)(Y_{1},\dots,Y_{k+l})=\frac{1}{(k+l)!}\mathrm{Sym}_{k+l}(f(Y_{1},\dots,Y_{k})g(Y_{k+1},\dots,Y_{k+l})).$
There is another deformation of the latter which is a simplest case of the Feigin–Odesskii shuffle algebras [12] (note that we use a different normalization constant).
@@
The rational Feigin–Odesskii shuffle algebra $\mathcal{S}^{\kappa}$ is the vector space
$\mathcal{S}^{\kappa}=\bigoplus\mathcal{S}^{\kappa}_{k},\qquad\mathcal{S}^{\kappa}_{k}=\mathbf{C}[Y_{1},\dots,Y_{k}]^{S_{k}}$
@@
$(f*^{\kappa}g)(Y_{1},\dots,Y_{k+l})$
$\qquad{}=\frac{1}{(k+l)!}\mathrm{Sym}_{k+l}\Biggl(\prod_{\begin{subarray}{c}i=1,\dots,k\\ j=k+1,\dots,k+l\end{subarray}}\frac{Y_{i}-Y_{j}+\kappa}{Y_{i}-Y_{j}}f(Y_{1},\dots,Y_{k})g(Y_{k+1},\dots,Y_{k+l})\Biggr).$
The $N$-truncated rational Feigin–Odesskii shuffle algebra $^{N}\!\mathcal{S}^{\kappa}$ is the vector space
$^{N}\!\mathcal{S}^{\kappa}=\bigoplus_{k=0}^{\infty}{}^{N}\!\mathcal{S}^{\kappa}_{k},\qquad{}^{N}\!\mathcal{S}^{\kappa}_{k}=\mathbf{C}[\hbar][Y_{1},\dots,Y_{k}]^{S_{k}}_{<N}$
@@
The original construction of elliptic algebras [12, 40, 41] defines a shuffle product with factor $\vartheta_{1}(z)$ (up to shifts) on the symmetric algebra $S^{\bullet}H^{0}(\Theta_{N})$ of global sections of a degree $N$ line bundle on an elliptic curve. By [10], the latter admits a degeneration such that $\vartheta_{1}(z)\mapsto z$} and the symmetric algebra becomes $^{N}\!\mathcal{S}^{\kappa}$.
The following theorem is a corollary of Proposition 3.7.
@@
One can similarly relate the trigonometric Shibukawa–Ueno operator to the Laurent polynomial representation of the non-degenerate affine Hecke algebra [6, Section 12.3] via the affinization procedure of [32]. It would be interesting to relate the elliptic Shibukawa–Ueno operators to elliptic affine Hecke algebras [21].
3.3 Shifted Yangian of $\boldsymbol{\smash{\mathfrak{sl}_{2}}}$
@@
The following definition is [17, Section 2.1]; we use its equivalent form in terms of currents. In particular, we assume that the weight $(m_{1},m_{2})$ is antidominant, i.e., $m_{1}\leq m_{2}$.
@@
The shifted Yangian $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})$ is an algebra over $\mathbf{C}[\hbar]$ generated by $\bigl\{x_{+}^{(i)},x_{-}^{(i)}\bigr\}_{i\geq 1}$ and $\bigl\{d_{j}^{(i)}\bigr\}_{i>m_{j}}$ for $j=1,2$ that we combine into the series
$x_{+}(u)=\sum_{i\geq 1}x_{+}^{(i)}u^{-i},\qquad x_{-}(u)=\sum_{i\geq 1}x_{-}^{(i)}u^{-i},\qquad d_{j}(u)=u^{-m_{j}}+\sum_{i>m_{j}}d_{j}^{(i)}u^{-i}$
with values in $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})\bigl(\bigl(u^{-1}\bigr)\bigr)$ satisfying
$[d_{1}(u),d_{2}(v)]=0,$ (3.8)
$(u-v)[x_{+}(u),x_{-}(v)]=\hbar\bigl(d_{1}(v)^{-1}d_{2}(v)-d_{1}(u)^{-1}d_{2}(u)\bigr),$ (3.9)
$(u-v)[d_{j}(u),x_{+}(v)]=\hbar(\delta_{j1}-\delta_{j2})d_{j}(u)(x_{+}(u)-x_{+}(v)),$ (3.10)
$(u-v)[d_{j}(u),x_{-}(v)]=\hbar(\delta_{j2}-\delta_{j1})(x_{-}(u)-x_{-}(v))d_{j}(u),$ (3.11)
$(u-v)[x_{\pm}(u),x_{\pm}(v)]=\mp\hbar(x_{\pm}(u)-x_{\pm}(v))^{2}.$ (3.12)
The positive ((negative)) subalgebra $\mathrm{Y}(\mathfrak{gl}_{2})^{\pm}$ is the subalgebra of $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})$ generated by the modes of $x_{\pm}(u)$. The non-negative subalgebra $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})^{\geq 0}$ (resp. non-positive subalgebra $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})^{\leq 0}$) is the subalgebra generated by the modes of $x_{+}(u)$, $d_{j}(u)$ (resp. $x_{-}(u)$, $d_{j}(u)$) for j=1,2{j=1,2}.
@@
$\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})^{\geq 0}\xrightarrow{\sim}\mathrm{Y}_{0,0}(\mathfrak{gl}_{2})^{\geq 0},\qquad x_{+}(u)\mapsto x_{+}(u),\qquad d_{j}(u)\mapsto u^{m_{j}}d_{j}(u),$
@@
Let $m=m_{1}-m_{2}$. The definition of the shifted Yangian $\mathrm{Y}_{m}(\mathfrak{sl}_{2})$ can be found in [4, Section B.1], see also its form in terms of currents $e(u)$, $f(u)$, $h(u)$ in [14, Section 6.1]. Its relation with $\mathrm{Y}_{m_{1},m_{2}}(\mathfrak{gl}_{2})$ is as follows.
Proposition 3.15 ([17, Proposition 2.19]).
There is a $\mathbf{C}[\hbar]$-algebra embedding
@@
Let $(m_{1},m_{2})=(-N,N)$. Let
$\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}=\mathbf{C}[\hbar][w_{1},\dots,w_{N}]\biggl[\frac{1}{w_{i}-w_{j}+k\hbar}\>\bigg|\>i,j=1,\dots,N,\,k\in\mathbb{Z}\biggr]$
be a localization of the algebra of functions on $\mathbb{A}^{N}\times\mathbb{A}^{1}$ (see also (5.4)). Denote by
$^{N}\!\tilde{\mathcal{A}}_{\hbar}:=\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}}\bigl[\mathbf{u}_{1}^{\pm},\dots,\mathbf{u}_{N}^{\pm}\bigr]$ (3.13)
the algebra of difference operators such that
$\mathbf{u}_{i}^{\pm}f(w)=f(w\pm\hbar\epsilon_{i})\mathbf{u}_{i}^{\pm},\qquad f(w)\in\mathrm{U}_{\hbar}(\mathfrak{h}_{N})^{\mathrm{gen}},$
where $f(w\pm\hbar\epsilon_{i}):=f(w_{1},\dots,w_{i}\pm\hbar,\dots,w_{N},\hbar)$. The homomorphism of the following theorem is called the GKLO homomorphism, see [20] for the non-shifted version, [29] for dominant shifts, and [4] for general definition. We use the version adapted to $\mathfrak{gl}_{2}$.
Theorem 3.16 ([17, Theorem 2.35]).
There is a $\mathbf{C}[\hbar]$-algebra homomorphism
@@
This map is not injective; for conjectural description of the kernel, see [4, Remark B.21]. The following is a special case of truncated shifted Yangians, see [4, Appendix B].
@@
Recall the shuffle algebra of Definition 3.8. The following is [45, Theorem 6.20] combined with Theorem 3.11, see also [11].
@@
4 Harish-Chandra bimodules
In this section, we recall the definition of the category of Harish-Chandra bimodules and the action of degenerate affine Hecke algebras on its certain objects. The descent of this action to the Kostant–Whittaker reduction of Section 6 will define a homomorphism from the shuffle algebra (and, more generally, truncated shifted Yangian of $\mathfrak{sl}_{2}$) to the quantum Toda lattice.